OpenStudy (anonymous):

if 3x+y\y = 6/5 what is the value of x/y

OpenStudy (accidentalaichan):

Well, the format of the question: "what is the value of (x/y)?" is weird, but here is how I would work it out: (3x+y)/y=6/5 (3x+y)=(6/5)y 3x=(1/5)y x=(1/15)y and there's your answer. :) hope this helped.

OpenStudy (anonymous):

what is the value of y ?

OpenStudy (anonymous):

can u explain in detail ??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

can u explain this step 3x=(1/5)y

OpenStudy (raden):

before that, is (3x+y)=(6/5)y subtract by y on both sides, we get 3x+y-y=(6/5)y - y 3x = (6/5)y - 5/5 y 3x = 1/5 y

OpenStudy (anonymous):

thanks

OpenStudy (raden):

welcome

OpenStudy (anonymous):

I have a doubt, how did this step come 3x = (6/5)y - (5/5) y ??